OpenStudy (anonymous):

How do I make a polynomial out of an equation that is Second degree; 2i (main question: how do I make a polynomial out of imaginary numbers?) Example of a solved one: Second dregree; 3 = x^2-6x+9

OpenStudy (anonymous):

start with \[(x-2i)(x+2i)\] then multiply

OpenStudy (anonymous):

get \[x^2+4\] pretty much in your head

OpenStudy (anonymous):

think "if the solution is 2i, then it must have been \[x^2=-4\]

jimthompson5910 (jim_thompson5910):

x = 2i or x = -2i .... note: complex solutions come in conjugate pairs x-2i = 0 or x+2i = 0 (x-2i)(x+2i) = 0 x^2-4i^2 = 0 x^2 - 4(-1) = 0 x^2 + 4 = 0 So the polynomial x^2 + 4 has the roots 2i and -2i

OpenStudy (anonymous):

Thanks.